The slope of a line is 20 and the point (3, 5) is on the line. Express the line in slope-intercept form.

y = 20x + -55


y = 3x + -55


y = 20x + -97


y = 3x + -35

Answer by nerdybill(7384)

The slope of a line is 20 and the point (3, 5) is on the line. Express the line in slope-intercept form.
.
Insert information into "point-slope" form of a line:
y - y1 = m(x - x1)
y - 5 = 20(x - 3)
y - 5 = 20x - 60
y - 5 = 20x - 60
y = 20x - 55
or
y = 20x + -55